# Unlock the Power of Frozen LLMs in Knowledge Graph Completion

### v2

@@ 3 Experiments
  • RQ1: Can frozen LLMs with the proposed probing methods effectively achieve performance on par with their fine-tuned counterparts?
@@
3.1 Datasets
Dataset |E| |R| # Train # Valid # Test
@@ 3.1 Datasets
Table 1: Statistical information of datasets.
@@
In our experiments, we employ five widely-used KG datasets for triple classification: FB13 (Socher et al., 2013), WN11 (Socher et al., 2013), FB15K-237N (Lv et al., 2022), WN18RR (Dettmers et al., 2018), and UMLS (Yao et al., 2019), as well as a commonly used KG dataset for relation prediction: YAGO3-10 (Dettmers et al., 2018).
@@
For FB13, we use Wikipedia entity descriptions as input sentences. For FB15K-237N, the entity descriptions provided by (Xie et al., 2016) are used. Synset definitions from (Yao et al., 2019) served as entity descriptions for WN18RR. In the case of UMLS, entity descriptions provided by the dataset are used. In the following, we refer to the entity descriptions as non-generated entity descriptions.
3.2 Experimental Settings
@@
To generate entity descriptions, for the first approach in Section 2.3, we concatenate all subgraph triples to generate descriptions. For the second approach, we employ GPT-3.5-turbo as the LLM. Due to cost considerations, we randomly select 4000 test examples from each dataset to generate entity descriptions. We mark the results of these methods with an asterisk (*). In the following text, we refer to these entity descriptions as generated entity descriptions (Tri)/(GPT).
@@
We compare our method with four structure information-based methods: TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), ComplEx (Trouillon et al., 2016) and RotatE (Sun et al., 2019), as well as seven additional information-based methods: KG-LLAMA (Yao et al., 2023), KG-BERT (Yao et al., 2019), KG-T5 (Saxena et al., 2022), LLaMA-7b, Alpaca-7B, LLaMA-7B-ICL and Alpaca-7B-ICL (Zhang et al., 2023). KG-LLAMA, KG-BERT and KGT5 fine-tune the language models using all the training samples. LLaMA-7b and Alpaca-7B using the frozen language model to predict. LLaMA-7B-ICL and Alpaca-7B-ICL using 2-shot samples with in-context learning to predict. Please refer to Section 4.1 for detailed introductions of these methods.
For structure information-based methods, we reproduce the results with OpenKE (Han et al., 2018). For additional information-based methods, we reproduce the results with official code implementation. For KGT5, there’s no public record or code for the method on the FB15K237N, WN18RR, UMLS dataset. Please refer to Appendix A.2 for detailed experimental settings.
Method Pred Samples FB13 WN11 15K237N WN18RR UMLS
TransE S-F ALL 0.815 0.759 0.697 0.884 0.845
DistMult S-F ALL 0.862 0.871 0.587 0.851 0.864
ComplEx S-F ALL 0.857 0.862 0.657 0.841 0.908
RotatE S-F ALL 0.819 0.847 0.685 0.882 0.921
KG-BERT L-MLP ALL 0.904 0.935 0.560 0.927 0.773
KGT5 N-L ALL 0.663 0.728 - - -
LLAMA-7B N-L 0 0.699 0.661 0.573 0.458 0.658
LLAMA-7B-ICL N-L 2 0.501 0.500 0.578 0.502 0.538
KG-LLAMA-7B N-L ALL 0.892 0.955 0.748 0.921 0.858
KG-LLAMA-SAMPLED N-L 3k/10k 0.727 0.682 0.614 0.574 0.834
LLAMA-MLP H-MLP 3k/10k 0.855 0.872 0.674 0.858 0.882
LLAMA-DES-H H-MLP 3k/10k 0.901 - 0.738 0.931 0.862
LLAMA-DES2 (Tri) H-MLP 3k/10k 0.842 0.755 0.671 0.779 0.782
LLAMA-DES2 (GPT)* H-MLP 3k/10k 0.912 0.917 0.726 0.924 0.860
Table 2: The main experiment results (accuracy) of triple classification. In the $Pred$ column, $S\mbox{-}F$ uses score function, $L\mbox{-}MLP$ uses MLP on last layer hidden states, $N\mbox{-}L$ uses natural language, and $H\mbox{-}MLP$ uses MLP on intermediate layer hidden states to predict. $Samples$ indicates training sample counts, with $3k/10k$ indicating $10k$ samples for the UMLS dataset and $3k$ for others. LLAMA-MLP uses LLaMA-7B with MLP for hidden state classification and $PT_{1}$ for stimuli generation; LLAMA-DES-H combines non-generated descriptions with $PT_{2}$ for stimulus generation; LLAMA-DES2 (Tri)/(GPT) uses first/second generated descriptions with $PT_{2}$ for stimuli generation. Top three results per dataset are bolded.
3.3 Main Results (For RQ1)
@@
From Table 2, it can be observed that LLAMA-DES-H and LLAMA-DES2 (GPT), when using a partial training set, exhibit comparable performance to KG-LLAMA-7B, the latter being fine-tuned with all training triples. Moreover, when trained on subsets of equivalent size, LLAMA-DES-H and LLAMA-DES2 (GPT) markedly surpass KG-LLAMA-SAMPLED, the existing state-of-the-art approach that relies exclusively on LLaMA for KGC. For more details, including the results of baseline models such as Alpaca-7B-ICL and Structural-aware-IT, and the performance of our method on the full training set, refer to Appendix B.4.
@@
The experimental results presented in Table 2 also validate the effectiveness of the approaches proposed in Section 2.2 and Section 2.3. Specifically, LLAMA-DES-H and LLAMA-DES2 (GPT) outperform LLAMA-MLP, demonstrating the effectiveness of using subgraph entity description generator to improve the LLM’s understanding of entities in the KG, as proposed in Section 2.3. Furthermore, LLAMA-MLP demonstrates superior performance compared to LLAMA-7B and LLAMA-7B-ICL, affirming the effectiveness of stimulation method proposed in Section 2.2. Thus, the experimental results substantiate the effectiveness of our proposed methodology. Additionally, the frozen LLM can also achieve similar KGC effects as that of fine-tuning the LLM.
@@
Notice that the performance on the UMLS dataset is special, where the use of generated descriptions hinders the model’s classification accuracy, and the use of non-generated entity descriptions is also worse than not using any entity descriptions. It is postulated that the underlying reason may lie in the highly specialized and domain-specific nature of entity names in the UMLS dataset, which are often precise medical terms with limited room for multiple interpretations. The entities in the knowledge graph have well-aligned with the internal entities of the LLM. Meanwhile, using entity descriptions may be affected by noisy description text, and descriptions generated by GPT-3.5-turbo may be influenced by hallucination, thereby affecting the classification effectiveness of the induced hidden states.
@@
In addition to the triple classification, we also validate the effectiveness of the proposed probing method on the relation prediction task as shown in Table 3. The experimental results similarly demonstrate that the probing method can achieve the nearly same performance as full training set fine-tuning, even when using only 0.6% of the training data. Detailed methodology and analysis are provided in Appendix B.1.
@@
Figure 3: Differences in KGC performance across different layers of the LLAMA-DES2 (GPT) [left] and Mistral-DES2 (GPT) [right]. The horizontal axis represents the position of the intermediate layers in LLaMA or Mistral (which have different backbones), and the vertical axis represents the dataset.
@@
LLAMA-DES2 (GPT)* H-MLP 6996 0.6824
Table 3: The experiment results (Hits@1) of relation prediction. LLAMA-MLP-7B refers to using LLaMA-7B as the base language model, using MLP as the hidden state classifier, and generating stimuli using $PT_{3}$; LLAMA-DES2 (GPT) refers to using the second generated description combined with $PT_{4}$ to generate stimuli.
3.4 Ablation Study (For RQ2)
Through the analysis depicted in the left portion of Figure 3, it is observed that with an increase in the index of the intermediate layer, the efficacy of hidden states for triple classification demonstrates an initial incline followed by a subsequent decline. Intermediate layers closer to natural language text output generally exhibit better prediction performance compared to the lowest layers. This observation corroborates the notion posited by Geva et al. (2020) that language models encode knowledge differentially across their Feed-Forward layers, with higher layers integrating information from lower layers. During pretraining, the lower layers may not have stored certain knowledge, and thus are unable to generate hidden states conducive to accurate triple classification. Furthermore, the prediction performance of the layers closer to the natural language text output is generally worse than those in the middle layers. This phenomenon may be attributed to the hallucination issue inherent in LLaMA, wherein higher layers could potentially either omit or alter triple information, leading to a blurred distinction between positive and negative triple hidden states induced by stimuli (Jiang et al., 2024; Zou et al., 2023). Detailed explanations can be found in Appendix B.5. Akin observations are discernible on the right side of Figure 3 for the Mistral model, which has a different backbone from LLaMA. Meanwhile, as observed in Table 5, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning with only a small sample size.
Regarding classification models, a comparative analysis encompassing MLP, SVM, and Logistic Regression, as delineated in Table 4, is conducted. The result shows that MLP outperforms all three datasets compared to the other two classification models, while Logistic Regression also exhibits results close to MLP. This underscores the efficacy of our stimulus methodology, whereby hidden states of positive and negative samples within a high-dimensional space are easily distinguishable.
@@ 3.4 Ablation Study (For RQ2)
Table 4: The variation in KGC prediction performance across different datasets using SVM, Logistic Regression, MLP as classification models. All these methods use LLaMA-7B as the base model, and generate stimuli using $PT_{1}$.
@@
Table 5: The variation in KGC prediction performance across different datasets using LLaMA, Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024) as base models (which have different backbones). Please refer to Table 2 for the number of training samples used by different methods.
3.5 Efficiency Study (For RQ3)
As illustrated in Figure 4, our method notably surpasses both KG-LLAMA and KG-BERT on the FB13 when trained with only 100 samples, as opposed to the 3000 samples used by the latter methods. For the UMLS, the performance of KG-LLAMA and KG-BERT with 3000 training samples is comparable to that of LLAMA-DES-H trained with only 500 samples. Additionally, combining the results from Table 2 and Table 10, for the FB13 dataset, our method attains 97.2% of the performance achievable with the full training set, utilizing only 500 samples, which constitutes a mere 0.079% of the total training samples. In contrast, KG-LLAMA and KG-BERT achieve only 78.7% and 63.1% of the maximum performance, respectively. For the UMLS, with 9.58% of the training set, our method reaches 90.6% of the full training set performance. For other datasets, please refer to Appendix B.6. Our experimental results confirm that the language model is capable of generating effective hidden states through the techniques described in Sections 2.2 and 2.3, demonstrating its potential for data-efficient scenarios within the context of KGC tasks.
Figure 4: The prediction performance of KG-LLAMA, KG-BERT, and LLAMA-DES-H on triple classification varies with the size of the training dataset. The horizontal axis represents the number of training samples used, with an equal number of positive and negative samples.
Meanwhile, we also evaluate the compute efficiency of our method. Table 8 shows that our method can achieve significantly lower consumption compared to parameter-efficient fine-tuning even in the setting of the full dataset. Compared to fine-tuned LLMs, we boost the GPU memory efficiency to 188×188$\times$188 × during model training and speed up 13.48×13.48$\times$13.48 × overall. Please refer to Appendix B.3 for details.
4 Related Work
4.1 Knowledge Graph Completion
Knowledge graphs are typically characterized by their inherent incompleteness, often containing substantial amounts of implicit or missing valuable knowledge (Hogan et al., 2021). The prevailing methodologies in KGC can be categorized into two main classes: structure information-based methods and additional information-based methods (Shen et al., 2022b). Structure information-based methods leverage the inherent structural data present within knowledge graphs. For example, RESCAL (Nickel et al., 2011), DistMult (Yang et al., 2014), COMPGCN (Vashishth et al., 2019), and ConvE (Dettmers et al., 2018) primarily use scoring functions to evaluate the semantic similarities of entity or relation embeddings within a latent embedding space to infer missing information. Additional information-based methods incorporate a variety of supplementary data to enrich the knowledge graphs. These methods utilize node attributes, entity-related information, or relational path data to enhance the features of knowledge embeddings critical for KGC tasks. The prominent examples are KG-BERT (Yao et al., 2019) and LASS (Shen et al., 2022a), which represent entities, relations, and triples as textual sequences, turning KGC into a sequence classification problem. KG-S2S (Chen et al., 2022) and KGT5 (Saxena et al., 2022) employ encoder-decoder architecture language models to complete KGC in a generative manner. KG-LLAMA (Yao et al., 2023) utilizes triples composed of entity and relation names as sentences, parameter-efficient fine-tuning the LLaMA model for KGC. Li et al. (2024) enhances KGC with LLM by generating entity descriptions and designing loss functions for fine-tuning.
4.2 Large Language Models
Several studies have explored the mechanisms through which Large Language Models process, interpret, and generate factual content. As highlighted by Saier et al. (2023), these models function as extensive repositories of knowledge, encapsulating a vast array of information about the world within their parameters. This insight into LLMs underscores their potential as powerful tools for a wide range of applications, including knowledge graph completion, by leveraging the rich informational content stored in their intricate network structures. Ramrakhiyani et al. (2023) have conducted evaluations across multiple Large Language Models to assess their knowledge of world geography using probing techniques. Similarly, Onoe et al. (2022) developed a methodology for automatically constructing cloze prompts by masking out spans, aiming to assess the language models’ understanding of entities through factual probing. In addition, Azaria and Mitchell (2023) and Zou et al. (2023) have explored training classifiers to estimate the truthfulness of statements based on the activations of the hidden layers in LLMs.
5 Conclusion
In this paper, we introduce a novel data-efficient probing method that completes KGC by leveraging the internal knowledge of LLMs. Experiments show that using the probing method with generated entity descriptions, frozen LLMs can achieve performance comparable to that of their fine-tuned counterparts. Simultaneously, this method saves a significant amount of GPU memory and time compared to parameter-efficient fine-tuning approaches. This work explores the potential of frozen LLMs in KGC tasks with both efficacy and efficiency. Compared to fine-tuning LLMs, we hit the core of KGC problem in a better way.
Limitations
Our current architectural design is only applicable to open-source LLMs or closed-source models that provide access to token embeddings; it is not suitable for closed-source models that exclusively output text. Recognizing these limitations, future work will aim to expand our architecture to accommodate closed-source LLMs that only output text or token probabilities. Additionally, future efforts could integrate structure information-based KG models, such as TransE, with our methodology in a hybrid approach.
Appendix A Experimental Details
A.1 Generate Entity Descriptions using GPT-3.5-turbo
In Section 2.3, we utilize LLM with the in-context learning technique to generate descriptions for entities based on the triples in their one-hop subgraph. In the experiments, we utilize the GPT-3.5 as the LLM. The prompt template used to generate entity descriptions is shown in Figure 5. Specifically, the $EN(e)$ is the entity name of entity $e$; the $\operatorname{concat\_subgraph()}$ is a function that first converts triples from the one-hop subgraph where the entity is located into sentences, and then concatenates these sentences using a specified delimiter. It is important to note that all the triples used in the prompt are from the training set, preventing potential data leakage from the test set in the generated entity descriptions. Table 6 shows the descriptions of entities generated using this method, thus achieving "entity alignment" between the knowledge graph and the LLM.
Entity Generated Description Non-Generated Description
Gustav Mahler
Gustav Mahler was a Jewish composer
known for his significant contributions
to the world of music. He lived in the
United States and was of German nationality.
Gustav Mahler (German; 7 July 1860 – 18 May 1911)
was an Austro-Bohemian late-Romantic composer,
and one of the leading conductors of his generation…
James Hayter
James Hayter was a male actor born in Scotland,
known for his profession in acting.
He passed away in Spain.
Henry James Goodenough Hayter
(23 April 1907 – 27 March 1983), better known
as James Hayter, was a British actor. He is best
remembered for his roles as Friar Tuck in the film The Story
of Robin Hood and His Merrie Men (1952) and…
Table 6: Comparison of generated entity descriptions with non-generated entity descriptions on the FB13 dataset.
Figure 5: The prompt template used to generate entity descriptions based on the one-hop subgraph where the entity is located for GPT-3.5-turbo.
A.2 Implementation Details
Figure 6: The prompt template used in Alpaca-7B-ICL on the WN18RR dataset.
In our implementation, we experimented with these baselines using the hyperparameters recommended in their paper.
For the negative sampling employed on the datasets, we utilize the same negative sampling technique as Yao et al. (2019) for FB13 and WN11. Specifically, FB15K-237N is designed to extract hard negative triples, providing a more challenging evaluation and reducing false negatives in the validation/test datasets during dataset construction. A similar approach is adopted for constructing negative samples for WN18RR and UMLS. Meanwhile, for all experiments in our study, the training sets contain an equal number of positive and negative samples.
For the MLP classifier configuration, we opt for a batch size of 64, a learning rate of 3e-5, and utilize the AdamW optimizer, with training restricted to 30 epochs. For LLaMA-7B-ICL and Alpaca-7B-ICL, we randomly select one positive sample and one negative sample from the training set of each dataset, respectively. The prompt for KGC using Alpaca-7B-ICL on the WN18RR dataset is illustrated in Figure 6. Additionally, the hyper-parameters for models trained with the LoRA technique, specifically KG-LLAMA-7B, are outlined in Table 7.
It’s worth noting that for Table 2, Table 3, Table 4, Table 5 and Figure 4, we determine the most suitable layer for the task by experimenting with the validation set, and then report the results on the test set. For Figure 3, the results shown are for different layers of the language model on the test set. For the baseline results in Table 2 and Table 3, we give priority to using the results reported in the official baseline paper. For results not included in the paper, we replicate the experiments using the hyperparameters recommended by the official sources and conduct multiple attempts and optimizations on these hyperparameters as well. Each experiment is executed three times with random initializations, and the mean results are reported.
Parameter Value
optimizer adamw
learning rate 3e-4
epoch 3
weight decay 0
LoRA modules q_proj, k_proj, v_proj, o_proj
LoRA r 8
LoRA alpha 16
LoRA dropout 0.05
Table 7: Hyper-parameters of KG-LLAMA-7B.
Appendix B Additional Results and Analysis
B.1 Relation Prediction
Figure 7: Prompt templates to stimulate the LLM for relation prediction.

### v3

@@ 3. Experiments
  • RQ1: Can LLMs with our method effectively achieve performance on par with their fine-tuned counterparts?
@@
3.1. Datasets
Dataset $|\mathbf{E}|$ $|\mathbf{R}|$ Train # Valid # Test
@@ 3.1. Datasets
Table 2. Statistical information of datasets.
@@
In our experiments, we employ five widely-used KG datasets for triple classification: FB13 (Socher et al., 2013), WN11 (Socher et al., 2013), FB15K-237N (Lv et al., 2022), WN18RR (Dettmers et al., 2018), and UMLS (Yao et al., 2019). For relation prediction, we employ the YAGO3-10 dataset (Dettmers et al., 2018), and for entity prediction, we use WN18RR for evaluation.
@@
For FB13, we utilize Wikipedia to construct entity descriptions. For FB15K-237N, the entity descriptions provided by (Xie et al., 2016) are used. Synset definitions from (Yao et al., 2019) served as entity descriptions for WN18RR. In the case of UMLS, entity descriptions provided by the dataset are used. In the following, we refer to these entity descriptions as non-generated entity descriptions.
3.2. Experimental Settings
@@
To generate entity descriptions, for the first approach in Section 2.3, we concatenate ten randomly selected triples from the one-hop subgraph surrounding each entity to construct the descriptions. For the second approach, we employ GPT-3.5-turbo as the LLM. In the following text, we refer to these entity descriptions as generated entity descriptions (Tri)/(GPT).
@@
We compare our method with four structure information-based methods: TransE (Bordes et al., 2013), DistMult (Yang et al., 2014), ComplEx (Trouillon et al., 2016) and RotatE (Sun et al., 2019), as well as seven additional information-based methods: KG-LLAMA (Yao et al., 2024), KG-BERT (Yao et al., 2019), KG-T5 (Saxena et al., 2022), LLaMA-7b, Alpaca-7B, LLaMA-7B-ICL and Alpaca-7B-ICL (Zhang et al., 2024). KG-LLAMA, KG-BERT and KGT5 fine-tune the language models using all the training samples. LLaMA-7B and Alpaca-7B rely on non-fine-tuned LLMs to directly perform predictions. LLaMA-7B-ICL and Alpaca-7B-ICL employ LLMs with in-context learning to make predictions (Zhang et al., 2024). We give priority to using the results reported in the official baseline paper. Please refer to Section 4.1 for detailed introductions of these baseline methods.
For the negative sampling employed on the datasets, we utilize the same negative sampling technique as (Yao et al., 2019). For the intermediate layer selection, we identify the most suitable layer for the task by experimenting with the validation set. Each experiment is executed three times with random initializations, and the mean results are reported.
3.3. Main Results (For RQ1)
@@
From Table 1, it can be observed that LLAMA-DES-H and LLAMA-DES2 (GPT), when using a partial training set, exhibit comparable performance to KG-LLAMA-7B, the latter being fine-tuned with all training triples. Moreover, when trained on subsets of equivalent size, LLAMA-DES-H and LLAMA-DES2 (GPT) markedly surpass KG-LLAMA-SAMPLED, the existing state-of-the-art approach that relies exclusively on LLaMA for KGC.
@@
The experimental results presented in Table 1 also validate the effectiveness of the approaches proposed in Section 2.2 and Section 2.3. Specifically, LLAMA-DES-H and LLAMA-DES2 (GPT) outperform LLAMA-MLP, demonstrating the effectiveness of using subgraph entity description generator to improve the LLM’s understanding of entities in the KG, as proposed in Section 2.3. Furthermore, LLAMA-MLP demonstrates superior performance compared to LLAMA-7B and LLAMA-7B-ICL, affirming the effectiveness of stimulation method proposed in Section 2.2. Thus, the experimental results substantiate the effectiveness of our proposed methodology. Additionally, the non-fine-tuned LLM can also achieve similar KGC effects as that of fine-tuning the LLM.
@@
Notice that the performance on the UMLS dataset is special, where the use of generated descriptions hinders the model’s classification accuracy, and the use of non-generated entity descriptions is also worse than not using any entity descriptions. It is postulated that the underlying reason may lie in the highly specialized and domain-specific nature of entity names in the UMLS dataset, which are often precise medical terms with limited room for multiple interpretations. The entities in the knowledge graph have well-aligned with the internal entities of the LLM. Meanwhile, using entity descriptions may be affected by noisy description text, and descriptions generated by GPT-3.5-turbo may be influenced by hallucination, thereby affecting the classification effectiveness of the induced hidden states. Meanwhile, our method shows slightly lower performance on the WN11 and FB15K237N datasets compared to KG-LLAMA, which fine-tuned on the full training set. This discrepancy may stem from the fact that the pretrained LLM did not acquire knowledge of certain triples in the KG during its pretraining phase, resulting in incorrect predictions even when probing is conducted.
@@
In addition to the triple classification, we also validate the effectiveness of the proposed probing method on the relation prediction task as shown in Table 3. The experimental results similarly demonstrate that the probing method can achieve the nearly same performance as full training set fine-tuning, even when using only 0.6% of the training data. For entity prediction, we obtain the plausibility confidence scores of the triples through the classifier and rank them accordingly. Table 4 also displays similar results, showing that after using model-friendly descriptions, the LLAMA-DES2 (GPT) method improves the performance of LLAMA-MLP-7B and surpasses its fine-tuned counterparts.
@@
Figure 3. Differences in KGC performance across different layers of the LLAMA-DES2 (GPT) [left] and Mistral-DES2 (GPT) [right]. The horizontal axis represents the position of the intermediate layers in LLaMA or Mistral, and the vertical axis represents the dataset.
@@
LLAMA-DES2 (GPT) H-MLP 6996 0.6824
LLAMA-DES2 (GPT) H-MLP ALL 0.7015
Table 3. The experiment results (Hits@1) of relation prediction.
Method Pred Samples WN18RR
KG-BERT L-MLP ALL 0.1102
StAR L-MLP ALL 0.2430
KGT5 N-L ALL 0.1011
KG-ChatGLM-6B N-L ALL 0.1613
LLaMA-7B N-L 0 0.0849
LLaMA-13B N-L 0 0.0991
KG-LLaMA-7B N-L ALL 0.2415
LLAMA-MLP-7B H-MLP 10k 0.2108
LLAMA-DES2 (GPT) H-MLP 10k 0.2433
LLAMA-DES2 (GPT) H-MLP ALL 0.2495
Table 4. The experiment results (Hits@1) of entity prediction.
3.4. Ablation Study (For RQ2)
Through the analysis depicted in the left portion of Figure 3, it is observed that with an increase in the index of the intermediate layer, the efficacy of hidden states for triple classification demonstrates an initial incline followed by a subsequent decline. Intermediate layers closer to natural language text output generally exhibit better prediction performance compared to the lowest layers. This observation corroborates the notion posited by (Geva et al., 2020) that language models encode knowledge differentially across their Feed-Forward layers, with higher layers integrating information from lower layers. During pretraining, the lower layers may not have stored certain knowledge, and thus are unable to generate hidden states conducive to accurate triple classification. Furthermore, the prediction performance of the layers closer to the natural language text output is generally worse than those in the middle layers. This phenomenon may be attributed to the hallucination issue inherent in LLaMA, wherein higher layers could potentially either omit or alter triple information, leading to a blurred distinction between positive and negative triple hidden states induced by stimuli (Jiang et al., 2024; Zou et al., 2023). Akin observations are discernible on the right side of Figure 3 for the Mistral model, which has a different backbone from LLaMA. This experimental results are consistent with the conclusions of (Jiang et al., 2024), which found that the MLP layers in the later stages have an inhibitory effect on the extracted knowledge, particularly contributing significantly to erroneous output decoding. As a result, the middle layers in their experiments also tend to outperform the later layers. Meanwhile, as observed in Table 6, our method demonstrates strong versatility across different backbone models, achieving performance comparable to full dataset fine-tuning with only a small sample size.
Regarding classification models, a comparative analysis encompassing MLP, SVM, and Logistic Regression, as delineated in Table 5, is conducted. The result shows that MLP outperforms all three datasets compared to the other two classification models, while Logistic Regression also exhibits results close to MLP. This underscores the efficacy of our stimulus methodology, whereby the model inherently encodes discriminative features for KGC tasks, making hidden states easily distinguishable in high-dimensional space and allowing straightforward classification boundaries to emerge without the need for complex, nonlinear decision functions.
@@ 3.4. Ablation Study (For RQ2)
Table 5. The variation in KGC prediction performance across different datasets using SVM, Logistic Regression, MLP as classification models. All these methods use LLaMA-7B as the base model, and generate stimuli using $PT_{1}$.
@@
Table 6. The variation in KGC prediction performance across different datasets using LLaMA, Mistral (Jiang et al., 2023), and Gemma (Team et al., 2024) as base models (which have different backbones). Please refer to Table 1 for the number of training samples used by different methods.